A particle of mass m starts from rest and moves in a circular path of radius r with uniform angular acceleration a. The kinetic energy of the particle after n revolutions is

Grade:10

1 Answers

deepak
7 years ago
by newtons eqn of motion, \omega ^2 = 2\alpha s
 
where \omega is angular velocity
s is distance covered (s=n2\pir)
and alpha is the angular acceleration(\alpha=a)
 
hence
\omega ^2 = 2a(4\pi n r)
we know that
v= \omega r      (where v is the linear velocity)
hence v2 = \omega^2 r^2
v2=n8a \pi r^3
since,
KE= ½ (m)v^2
=>K.E.=1/2×(m)×n8a \pi r^3
 
 
hope this is useful and p.s. dont do these kinds of problems involving rotational motion.
do it only after eleventh and take it slow and steady.
XD